Thank you for helping with that last question. I have another.
Please help!
15-w/4=28
Answer is -52

Answer by stanbon(75887)

15-w/4=28
---
Subtract 15 from both sides to get:
-w/4 = 13
Multiply both sides by -4 to get:
w = -4*13
w = - 52
================
Cheers,
Stan H.